# Representative Action Selection for Large Action Space Meta-Bandits

### v3

@@ 2 Subset Selection Framework
$\boxed{\mu_{a}(\theta):=\langle a,\theta\rangle,\quad\forall a\in\mathcal{A}_{\operatorname*{full}},\theta\sim\mathcal{N}(0,I).}$ (3)
We define the extreme points as those $x\in\mathcal{A}_{\operatorname*{full}}$ for which there do not exist distinct $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$ and $\lambda\in(0,1)$ such that $x=\lambda a+(1-\lambda)a^{\prime}$. By the extreme point theorem, if we select all extreme points—denoted $\mathcal{A}=\{a_{1},\dots,a_{K}\}$—as representatives of the full action space, the regret is zero. This is because any $a\in\mathcal{A}_{\operatorname*{full}}$ can be expressed as a convex combination of the extreme points: $a=\lambda_{1}a_{1}+\dots+\lambda_{K}a_{K}$, where $\lambda_{i}\geq 0$ and $\sum_{i=1}^{K}\lambda_{i}=1$. Thus, for any $\theta\in\mathbb{R}^{n}$:
$\begin{split}\langle a,\theta\rangle=\sum_{i=1}^{K}\lambda_{i}\langle a_{i},\theta\rangle\leq\sum_{i=1}^{K}\lambda_{i}\max_{a^{\prime}\in\mathcal{A}}\langle a^{\prime},\theta\rangle=\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}.\end{split}$
By equation 3, it yields $\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\leq\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}\leq\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}},$ where the left inequality uses $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$. Thus, the two quantities $\max_{a\in\mathcal{A}}\mu_{a}$ and $\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}$ are equal.
This example highlights how a geometric approach can be used to solve the stochastic combination problem. However, even if one only needs the extreme points, the set of extreme points may still be large, e.g, the extreme points of a Euclidean ball is infinite. To address this, we will later introduce the notions of $\epsilon$-nets. Without loss of generality, we assume $\mathcal{A}_{\operatorname*{full}}$ consists only of the extreme points of $\mathcal{A}_{\operatorname*{full}}$, as they are the only points of interest.
@@ 2.1 Epsilon Nets
If $\mathcal{A}_{\operatorname*{full}}$ is very large, a natural approach is to construct a grid over the action space, where the grid points serve as representative actions. This ensures that for every action in the full space, there exists a representative that is close to it. This idea is formally captured by the notion of a (geometric) $\epsilon$-net.
To proceed, we clarify what we mean by an $\epsilon$-net, as there are at least two definitions: one from a geometric perspective (vershynin2018high, , Chapter 4) and another from a measure-theoretic perspective (matousek2013lectures, , Chapter 10). Let $\|\cdot\|_{2}$ denote the Euclidean norm. Define the diameter of a compact set $r\in\mathbb{R}^{n}$ as $\mathrm{diam}(r):=\max_{a,b\in r}\|a-b\|_{2}$.
@@
  • Let $\mathcal{R}$ be a finite partition of the extreme points into disjoint clusters such that $\cup_{r\in\mathcal{R}}\;r=\mathcal{A}_{\operatorname*{full}}$. Given a measure $q$ assigning a value to each cluster $r\in\mathcal{R}$. A subset $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$ is called a Measure-Theoretic $\epsilon$-net with respect to measure $q$ if, for any cluster $r\in\mathcal{R}$, we have:
@@
A geometric $\epsilon$-net ensures small regret because if two actions $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$ are close in the Euclidean sense, then the deviation between $\mu_{a}$ and $\mu_{a^{\prime}}$ is small in the $L^{2}$-sense (i.e., their expected squared difference is small): $\|\mu_{a}\!-\!\mu_{a^{\prime}}\|_{L^{2}}=\!\left(\mathbb{E}(a\!-\!a^{\prime})^{\top}\theta\theta^{\top}(a\!-\!a^{\prime})\right)^{1/2}\!=\|a\!-\!a^{\prime}\|_{2},$ where $\theta^{\top}$ denotes the transpose of $\theta$. The equalities use equation 3 and $\mathbb{E}\theta\theta^{\top}=I$. Therefore, by definition, a geometric $\epsilon$-net guarantees the existence of an action $a\in\mathcal{A}$ whose expected outcome $\mu_{a}$ is close to that of the optimal action for any given bandit instance. However, this net suffers from the curse of dimension: e.g., for $0,1]^{n}$, the number of points needed to form a geometric $\epsilon$-net grows as $(1/\epsilon)^{n}$.
The measure-theoretic $\epsilon$-net addresses this issue. Put simply, the measure-theoretic $\epsilon$-net restricts the grid construction to only the most important clusters $r\in\mathcal{R}$, as determined by the $q$-measure.
2.2 Epsilon Net Algorithm
We propose Algorithm 1, a variant of the $\epsilon$-net algorithm originally introduced by haussler1986epsilon . It selects $K$ i.i.d. random actions, aligned with the distribution of bandit instances. Since repetitions are allowed, the resulting subset $\mathcal{A}$ may have fewer than $K$ distinct actions.
Algorithm 1 Epsilon Net Algorithm
2:Output: A subset of actions $\mathcal{A}$.
[…]
4:for $1,\dots,K$ do
5:  Sample a bandit instance $\theta$.
6:  Find optimal action $a^{*}(\theta)\!:=\!\operatorname*{arg\,max}_{a\in\mathcal{A}_{\operatorname*{full}}}\langle a,\theta\rangle$.
7:  $\mathcal{A}\leftarrow\mathcal{A}\cup\{a^{*}\}$ $\triangleright$ Repetition of actions is allowed
8:end for
We define the optimal action in a bandit instance $\theta$ as
$a^{*}(\theta):=\operatorname*{arg\,max}_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta).$
Assumption 2 (Unique optimal action).
The optimal action $a^{*}(\theta)$ is unique with probability $1$ over all bandit instances.
Define the Importance Measure $q$ over a partition $\mathcal{R}$:
$\boxed{q(r):=\Pr[a^{*}(\theta)\in r]=\!\int\!\mathbb{1}\{a^{*}(\theta)\in r\}\;p(\theta)d\theta,}$ (4)
where $p(\theta)$ is the density of $\theta$ and $\int p(\theta)d\theta=1$. Under Assumption 2, measure $q$ is a probability distribution. It reflects the probability that a given cluster contains the optimal action and thus represents the potential contribution of that cluster to the expected regret.
Assumption 3.
The support of measure $q$ is compact.
By definition, Algorithm 1 samples $K$ i.i.d. extreme points from clusters in $\mathcal{R}$ according to measure $q$. If a cluster $r\in\mathcal{R}$ has a higher measure $q(r)$, its elements are more likely to be included in the output. In fact, with high probability, this algorithm outputs a measure-theoretic $\epsilon$-net of $\mathcal{A}_{\operatorname*{full}}$ with respect to measure $q$. (This is a simplified version of Theorem 10.2.4 of matousek2013lectures .)
Lemma 4.
Given a partition $\mathcal{R}$ of the full action space, and the importance measure $q$ assigning a value to each cluster $r\in\mathcal{R}$. Let $\mathcal{A}$ be the output of Algorithm 1 after $K$ samples. Then, with probability at least $1-\frac{1}{\epsilon}\exp(-K\epsilon)$, it holds that for any cluster $r\in\mathcal{R}$,
@@ 2.2 Epsilon Net Algorithm
The partition $\mathcal{R}$ bridges the two definitions of $\epsilon$-net: choosing one point from each cluster gives an $\epsilon$-net in both senses, though with different values of $\epsilon$. Geometrically, $\epsilon$ is the largest cluster diameter $\epsilon:=\max_{r\in\mathcal{R}}\mathrm{diam}(r)$; measure-theoretically, $\epsilon$ is the smallest cluster measure $\epsilon:=\min_{r\in\mathcal{R}}q(r)$.
@@ 3 Regret Analysis
In this section, we begin by analyzing a special class of geometric $\epsilon$-nets, constructed by partitioning the action space into clusters and selecting a single representative action from each cluster. We then extend the analysis to obtain algorithm-dependent bounds for the output of Algorithm 1.
Definition 1 (Reference subsets).
Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. A reference subset is a set $\mathcal{A}:=\{a_{1},\dots,a_{m}\}$, where each representative $a_{\ell}\in\mathcal{A}$ corresponds to a cluster $r_{\ell}$ and each cluster $r_{\ell}$ is contained within a closed Euclidean ball of radius $\epsilon$ centered at $a_{\ell}$, i.e., $r_{\ell}\subset B(a_{\ell},\epsilon)$.
@@
Only for lower bounds, we assume a well-separated partition: if the optimal action lies in cluster $r$, then all actions in $r$ outperform those outside it.
Assumption 5.
@@
Theorem 6 (Regret bounds of reference subsets).
Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the full action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$, and an arbitrary reference subset $\mathcal{A}$. Then, there is an absolute constant $C>0$, such that
@@
$\mathbb{E}_{\theta}[\operatorname*{Regret}]\leq\max_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]+C\epsilon\sqrt{\log|\mathcal{R}|}.$
If the partition $\mathcal{R}$ satisfies Assumption 5, then
$\mathbb{E}_{\theta}[\operatorname*{Regret}]\geq\min_{r\in\mathcal{R}}\mathbb{E}_{\theta}\left[\max_{a\in r}\mu_{a}\right]-C\epsilon\sqrt{\log|\mathcal{R}|}.$
Proof sketch. For each cluster $\ell\leq m$, define a simple Gaussian process $\left\{Z_{a}\right\}_{a\in r_{\ell}}$ where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$. Define a non-negative random variable $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}$. When $a^{*}(\theta)\in r_{\ell}$, regret is upper bounded by $Y_{\ell}$ (or equal to it under Assumption 5). Thus, for any bandit $\theta$, the regret is bounded between $\min_{\ell\leq m}Y_{\ell}$ and $\max_{\ell\leq m}Y_{\ell}$. Finally, the expectations $\mathbb{E}\left[\min_{\ell\leq m}Y_{\ell}\right]$ and $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]$ can be bounded via concentration property of Gaussian process. $\square$
For each $\ell\leq m$, the term $\mathbb{E}_{\theta}[\max_{a\in r_{\ell}}\mu_{a}]$ equals $\mathbb{E}_{\theta}[\max_{a\in r_{\ell}}\mu_{a}-\mu_{a_{\ell}}]$, since $\mathbb{E}_{\theta}[\mu_{a}]=0$.
3.1 Regret Bounds of Algorithm
The algorithm’s regret bound is established by comparing it to that of a reference subset, for which we already have known expected regret bounds. The key difference is that whereas the reference subset includes a representative from each cluster, the algorithm may miss some clusters. However, the algorithm still achieves regret comparable to that of the reference subset, as it tends to miss clusters that contribute minimally to the expected regret.
The expected regret in previous results is taken over bandit instances $\theta$. In contrast, since the output of Algorithm 1 is random, the expected regret analyzed in this section is taken with respect to both the algorithm’s randomness (i.e., the sampled $\mathcal{A}$) and the distribution over $\theta$.
Theorem 7.
Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of the action space, with $\epsilon:=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$. Let $\mathcal{A}$ be output of Algorithm 1. For the same constant $C>0$ in Theorem 6,

### v4

@@ 3.1 Gaussian Processes
$\mu_{a}(\theta):=\langle a,\theta\rangle,\quad\forall a\in\mathcal{A}_{\operatorname*{full}},\theta\sim\mathcal{N}(0,I).$} (5)
Here, $\left$$\cdot$$\cdot\right$> is the standard inner product on $\mathbb{R}^{n}$, and $\mathcal{N}(0,I)$ is the standard Gaussian distribution on $\mathbb{R}^{n}$. Thus, in terms of the notation in Section 1, the family of bandits $\mathcal{F}$ may be identified as $\mathbb{R}^{n}$, and the distribution $\mathcal{P}$ as the standard Gaussian. Clearly, the family of variables $\left\{\mu_{a}\right\}_{a\in\mathcal{A}_{\operatorname*{full}}}$ given by (5) is a Gaussian process. Conversely, we note that any finite dimensional Gaussian process can be represented as (5), and any infinite dimensional Gaussian process [^1] Under very mild technical assumptions can be arbitrarily approximated by (5), in a sufficiently high dimension $n$. See (Vershynin, 2018, Chapter 7). Thus, we may assume the form (5) without loss of generality.
Moreover, crucially, the dimension $n$ will not be important in our bounds, and can be large. On the other hand, however, it may be also of interest to note that when $n$ is fixed, the setup (5) is equivalent to the Bayesian linear bandit setup, (Lattimore & Szepesvári, 2020) (although we solve a different problem). The relation with linear bandits will be discussed further in Remark 4.8.
We define the $L_{2}$ distance induced on the $\mathcal{A}_{\operatorname*{full}}$ by the process to be the distance
$\left\|\mu_{a}-\mu_{a^{\prime}}\right\|^{2}_{L_{2}}:=\mathbb{E}_{\theta}\left|\left<a,\theta\right>-\left<a^{\prime},\theta\right>\right|^{2}.$ (6)
Note that under the definition (5), the $L_{2}$ distance and the standard Euclidean distance in $\mathbb{R}^{n}$ coincide, $\left\|\mu_{a}-\mu_{a^{\prime}}\right\|^{2}_{L_{2}}=\left\|a-a^{\prime}\right\|_{2}^{2}$ for all $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$.
@@ 3.2 Epsilon Nets
There will be two senses in which we would be interested in approximating $\mathcal{A}_{\operatorname*{full}}$: First, for every $a\in\mathcal{A}_{\operatorname*{full}}$ we would like to have $a^{\prime}\in\mathcal{A}$ such that $a^{\prime}$ is close to $a$ in the distance (6) (or, equivalently, in $\left\|\cdot\right\|_{2}$). Second, in order for the subset selection algorithm to able to discover the points $a^{\prime}$, we need an additional, probabilistic notion of approximation.
The following two notions of an $\epsilon$-net, the geometric and the measure theoretic nets, formalize these approximations. We refer to (Vershynin, 2018, Chapter 4) for additional details on geometric nets, and to (Matousek, 2013, Chapter 10) for measure theoretic nets.
Definition 3.1.
@@
As discussed earlier, geometric $\epsilon$-nets will be used as approximations of the set $\mathcal{A}_{\operatorname*{full}}$. While we’ll assume that good approximating nets exist in the data, these nets will never need to be explicitly constructed. Instead, part of such representative points will be implicitly chosen by the subset selection algorithm.
The $\epsilon$-covering number of the set $\mathcal{A}_{\operatorname*{full}}$, denoted $N(\mathcal{A}_{\operatorname*{full}},\epsilon)$, is the size of the smallest $\epsilon$-net of $\mathcal{A}_{\operatorname*{full}}$:
[…]
Now, Let $\mathcal{R}$ be a finite partition of $\mathcal{A}_{\operatorname*{full}}$ into disjoint clusters $r$, such that $\cup_{r\in\mathcal{R}}\;r=\mathcal{A}_{\operatorname*{full}}$. Let $q$ be a probability distribution on $\mathcal{A}_{\operatorname*{full}}$, which therefore assigns a weight $q(r)$ to each cluster $r\in\mathcal{R}$.
Definition 3.2.
A subset $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$ is called a measure-theoretic $\epsilon$-net with respect to measure $q$ and partition $\mathcal{R}$, if, for any cluster $r\in\mathcal{R}$, we have:
@@
The measure theoretic $\epsilon$-net ensures that in a given partition $\mathcal{R}$, either $\mathcal{A}$ intersects a cluster $r$, or the cluster $r$ is small in terms of the distribution $q$. Roughly speaking, this will be used to show that if the subset selection algorithm doesn’t find a cluster $r$ of actions, then the cluster’s contribution to the total expected regret is small and can be ignored.
An important property of random sampling is that with high probability it will produce a measure theoretic $\epsilon$-net for any fixed partition of the support of the distribution. Specifically, we have the following result:
Lemma 3.3.
Fix $\epsilon>0$. Let $\mathcal{R}$ be a partition of $\mathcal{A}_{\operatorname*{full}}$, and let $q$ be an importance measure given by (4). Let $\mathcal{A}$ be the output of Algorithm 1 after $K$ samples. Then, with probability at least $1-\frac{1}{\epsilon}\exp(-K\epsilon)$, it holds that for any cluster $r\in\mathcal{R}$,
@@
This result is a special case of Theorem 10.2.4 in (Matousek, 2013), which in turn follows the analysis in (Haussler & Welzl, 1986). The algorithm in (Haussler & Welzl, 1986) is called the $\epsilon$-net Algorithm, and we have retained that name. The full proof of the Lemma is given in Supplementary D.
@@ 4 Regret Analysis
In this Section we introduce and discuss the regret bounds for subset selection Algorithm 1. Our main result is Theorem 4.5, which bounds the regret in terms of the covering numbers of the set $\mathcal{A}_{\operatorname*{full}}$. Typically, sets which exhibit clustering, or sets with low dimensional manifold structure, would have low covering numbers, and we expect such sets to commonly occur in practical problems (the manifold hypothesis, (Bengio et al., 2013)). Roughly speaking, suppose that the set $\mathcal{A}_{\operatorname*{full}}$ can be represented as a union of a few small clusters, as illustrated in Figure 1. The main term in regret bound bound in Theorem 4.5 is related to the complexity of each individual cluster, and typically would controlled by its diameter $\epsilon$. Additional terms involve a logarithmic dependence on the total number of clusters (times $\epsilon$), and an error term which accounts for the probability of Algorithm 1 to find the relevant clusters. This term decays exponentially with $K$.
Figure 1: Illustration of Theorem 4.5. The full action space (blue dots) lies on the unit sphere in $\mathbb{R}^{3}$ and can be partitioned into three clusters. The regret bound established in the theorem depends not on the diameter of the entire action space—indicated by the red dashed line—but rather on the diameters of the individual clusters, represented by the green dashed lines, each of which is significantly smaller.
Perhaps the most central technical notion in our analysis is the notion of partition $\mathcal{R}=\left\{r_{\ell}\right\}_{\ell\leq m}$, such that $\mathcal{A}_{\operatorname*{full}}=\cup_{l\leq m}r_{\ell}$, and we will typically be interested in partitions where each subset $r_{\ell}$ also have a relatively small diameter $\epsilon$. There are two main ways in which such partitions may be present in $\mathcal{A}_{\operatorname*{full}}$. First, some realistic datasets may be naturally clustered, as depicted in Figure 1. Typically, in such situations there will be a characteristic scale $\epsilon$, such that there are few clusters with diameter $\epsilon$, which are well separated. However, inside each cluster, on scales smaller then $\epsilon$, there may be no additional structure, i.e. the clusters themselves would look dense and high dimensional. In such situations, we still may benefit from a regret of order $\epsilon$. The second scenario is when the whole set $\mathcal{A}_{\operatorname*{full}}$ is a low dimensional manifold, typically in a high dimensional space. In that case, for any scale $\epsilon$ we construct relatively small $\epsilon$-nets, which can be used to derive appropriate partitions.
Section 4.1 introduces bounds on the regret in terms of the partitions and their diameters. These are the bounds that one could obtain for a fixed a-priori known partitions. In Section 4.2 we obtain bounds for the subset selected by Algorithm 1, which finds (an approximation of) the reference points for the best partition available in $\mathcal{A}_{\operatorname*{full}}$, without knowing this partition. Theorem 4.5 provides bounds in terms of a general partition given scale $\epsilon$, while in Theorem 4.9, by taking $\epsilon$ to 0 we provide worst case diminishing bounds.
4.1 Partitions and Reference Set Bounds
For a point $a\in\mathbb{R}^{n}$ and scalar $\epsilon\geq 0$ denote by $B(a,\epsilon):=\left\{a^{\prime}\in\mathbb{R}^{n}\hskip 5.69054pt|\hskip 5.69054pt\left\|a-a^{\prime}\right\|_{2}\leq\epsilon\right\}$ the closed Euclidean ball of radius $\epsilon$ around $a$.
Definition 4.1 ($\epsilon$-reference sets).
Consider a partition $\mathcal{R}:=\{r_{\ell}\}_{\ell\leq m}$ of $\mathcal{A}_{\operatorname*{full}}$. For an $\epsilon>0$, a set $\mathcal{A}^{\epsilon}:=\{a_{1},\dots,a_{m}\}$ is an $\epsilon$-reference set for $\mathcal{R}$ if for every $\ell\leq m$, $r_{\ell}\subseteq B(a_{\ell},\epsilon)$. The point $a_{\ell}$ is called the reference point of the subset $r_{\ell}$.
Note that the $\epsilon$-reference set is in particular a (geometric) $\epsilon$-net of $\mathcal{A}_{\operatorname*{full}}$. However, for our purposes it will be more convenient to discuss partitions, which we do with the notion of reference sets. Note also that given a partition $\mathcal{R}$, one can always find a reference set with $\epsilon=\max_{r\in\mathcal{R}}\operatorname{diam}(r)$.
Theorem 4.2 (Regret bounds of $\epsilon$-reference subsets).
Let $\mathcal{R}$ be a partition of $\mathcal{A}_{\operatorname*{full}}$ and let $\mathcal{A}^{\epsilon}$ be an $\epsilon$-reference subset for $\mathcal{R}$. There is an absolute constant $C>0$, such that
$\mathbb{E}_{\theta}[\operatorname*{Regret}]\leq\max_{\ell\leq m}\mathbb{E}_{\theta}\left[\max_{a\in r_{\ell}}\left(\mu_{a}-\mu_{a_{\ell}}\right)\right]+C\epsilon\sqrt{\log m}.$
Proof sketch. For each cluster $\ell\leq m$, define a simple GP $\left\{Z_{a}\right\}_{a\in r_{\ell}}$ where $Z_{a}:=\mu_{a}-\mu_{a_{\ell}}$. Define a non-negative random variable $Y_{\ell}:=\sup_{a\in r_{\ell}}Z_{a}$. When $a^{*}(\theta)\in r_{\ell}$, regret is upper bounded by $Y_{\ell}$. Thus, for any bandit $\theta$, the regret is bounded between $\min_{\ell\leq m}Y_{\ell}$ and $\max_{\ell\leq m}Y_{\ell}$. Finally, the expectations $\mathbb{E}\left[\min_{\ell\leq m}Y_{\ell}\right]$ and $\mathbb{E}\left[\max_{\ell\leq m}Y_{\ell}\right]$ can be bounded via appropriate concentration arguments. $\square$
@@ 4.1 Partitions and Reference Set Bounds
For the lower bounds, we’ll require an additional separation assumption: if the optimal action lies in cluster $r$, then all actions in $r$ outperform those outside it.
Assumption 4.3.
@@
Theorem 4.4.
If the partition $\mathcal{R}$ satisfies Assumption 4.3, then
@@
$\mathbb{E}_{\theta}[\operatorname*{Regret}]\geq\min_{\ell\leq m}\mathbb{E}_{\theta}\left[\max_{a\in r_{\ell}}\left(\mu_{a}-\mu_{a_{\ell}}\right)\right]-C\epsilon\sqrt{\log m}.$
4.2 Regret Bounds for the Algorithm
As discussed above, Theorem 4.5 provides regret bounds for the regret of the set $\mathcal{A}$ selected by Algorithm 1.
To get an intuition about the argument, suppose that $\mathcal{A}_{\operatorname*{full}}$ admits a good partition $\mathcal{R}$ (i.e. relatively few clusters, of small diameter $\epsilon$). We then show that for any fixed partition $\mathcal{R}$, samples in Algorithm 1 will hit most of the clusters of that partition, and can consequently be used as reference set points in Theorem 4.2. Indeed, by Lemma 3.3, either Algorithm 1 will hit the cluster, or the probability of that cluster under $q$ is small, in which case its contribution to regret will be small too. This gives rise to the additional error term in Theorem 4.5 compared to Theorem 4.2, but this term decays exponentially with $K$.
Theorem 4.5.
Let $\mathcal{R}$ be a partition of $\mathcal{A}_{\operatorname*{full}}$ and set $\epsilon=\max_{r\in\mathcal{R}}diam(r)$. Let $\mathcal{A}$ be the output of Algorithm 1 after $K$ iterations. Then